OpenStudy (anonymous):

solve the equation for x. write the exact solution and the approximate solution to the nearest hundredth, when appropriate. ln(x-5)=ln(3x+1)

OpenStudy (anonymous):

If you have written this problem correctly, you may want to go back and check. There is no reall number value of x that will satisfy this equation.

OpenStudy (anonymous):

You could set it in your calculator as y1 = ln(x-5) and Y2 = (3x+1) and you will find that they don't equal the same x using real numbers. But algebraically you can solve it to get x=-3

OpenStudy (zehanz):

If ln(x-5)=ln(3x+1), then x-5=3x+1, which means: -5=2x+1 and so 2x=-6, or x=-3. Seems OK, or? Let's put it back into the original equation: ln(-3-5) = ln(-8) ln(3*-3+1)=ln(-8). Although it looks ok, there is a problem: the ln-function cannot be calculated for negative numbers (or even 0)! So no real solutions exist.

OpenStudy (anonymous):

Took the words right outta my...ummm...fingers @ZeHanz

OpenStudy (zehanz):

It is important to check what kind of numbers are allowed beforehand, so as to avoid dissapointment :D

OpenStudy (zehanz):

@calmat01: sorry about that!

OpenStudy (anonymous):

lol...np!

OpenStudy (anonymous):

thank you everyone! @ZeHanz @TRowstudy @calmat01

OpenStudy (zehanz):

YW!